if f(x+1/x)=x^2+1/x^2, find f(x^3) please help me with this.

Grade:12

1 Answers

suvidhi mehta
6 years ago
we can write above equation as 
f(x+1/x) = (x+1/x)^2 – 2
now put (x+1/x) = y
so f(y) = y^2 -2
so f(x^3) = x^6 -2